# Thermodynamic and energetic constraints on out-of-equilibrium tunneling rates

### v1

@@ II Theoretical model
We study a bipartite system with Hamiltonian $\hat{H}_{0}=\hat{H}_{\text{L}}+\hat{H}_{\text{R}}$, where subsystems L (left) and R (right) may be taken as generic systems, possibly with strong interactions. They are coupled to each other by the tunneling Hamiltonian
@@
which we assume to be a small perturbation, and which induces transitions in the subsystems. Further, in the weak tunnel-coupling regime, we describe the total system with the product state $\hat{\rho}=\hat{\tau}_{\text{L}}\otimes\hat{\tau}_{\text{R}}$, where $\hat{\tau}_{\alpha}\equiv\exp({-\beta_{\alpha}\hat{H}_{\alpha}})/Z_{\alpha}$ is the Gibbs state at inverse temperature $\beta_{\alpha}$ with partition function $Z_{\alpha}\equiv\text{Tr}_{\alpha}\left\{\exp{(-\beta_{\alpha}\hat{H}_{\alpha}})\right\}$. With this, we express the rates $\Gamma_{\rightleftarrows}$, induced by the coupling, to absorb or emit a quantum of energy $\hbar\omega$ in the spectral representation [17], as
@@
$\begin{split}\Gamma_{\rightarrow}&\equiv\frac{2\pi}{\hbar}\sum_{nmlk}|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon^{\text{(L)}}_{mn}+\epsilon^{\text{(R)}}_{kl}-\hbar\omega)p^{\text{(L)}}_{n}p^{\text{(R)}}_{l},\\ \Gamma_{\leftarrow}&\equiv\frac{2\pi}{\hbar}\sum_{nmlk}|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon^{\text{(L)}}_{mn}+\epsilon^{\text{(R)}}_{kl}-\hbar\omega)p^{\text{(L)}}_{m}p^{\text{(R)}}_{k},\end{split}$ (6)
with their dependence on $\omega$ and inverse temperatures $\beta_{\alpha}$ left implicit. Here, $\mathcal{A}_{nmlk}\equiv\braket{mk}{\hat{A}}{nl}$ is the matrix element of $\hat{A}$ in the basis of the eigenstates of the generic Hamiltonian $\hat{H}_{0}$, namely $\hat{H}_{\text{L}}\ket{nl}=\epsilon^{\text{(L)}}_{n}\ket{nl},$ and $\hat{H}_{\text{R}}\ket{nl}=\epsilon^{\text{(R)}}_{l}\ket{nl}$. Furthermore, we have defined the energy differences $\epsilon^{(\alpha)}_{mn}\equiv\epsilon^{(\alpha)}_{m}-\epsilon^{(\alpha)}_{n}$, and the occupation probability of the state |$\ket{n}$ as $p^{(\alpha)}_{n}\ket{n}=\hat{\tau}_{\alpha}\ket{n}$. Note that the partition of the system into two subsystems is not required to derive Eq. (6) (see Appendices A, B and references therein) but is needed to implement a well-defined temperature bias. The scope of the present paper is to find bounds for these rates in the presence of such a temperature bias. It is therefore important to point out here that the temperature dependence of the tunneling rates enters only through the occupation probabilities $p_{n}^{(\alpha)}$.
The two tunneling rates in Eq. (6) are the key quantities that determine the transitions between subsystems L and R. To show this connection, we consider an observable, represented by the operator $\hat{Q}$ that commutes with $\hat{H}_{0}$, i.e. $[\hat{Q},\hat{H}_{0}]=0$, but does not commute with the tunneling Hamiltonian $\hat{V}(t)$. For simplicity, we here focus on operators satisfying $[\hat{Q},\hat{A}]=q\hat{A}$, meaning that the transitions induced by the operator $\hat{A}$ transfer exactly one quantum $q$. Nonetheless, more general results that do not rely on this assumption are provided in Appendix B. The rate of change of the observable $\hat{Q}$ defines the current operator
$\hat{I}(t)=-\frac{iq}{\hbar}\left(\hat{A}e^{-i\omega t}-\hat{A}^{\dagger}e^{i\omega t}\right).$ (7)
[…]
In terms of the rates (6), the transport quantities $I$ and $\mathcal{S}$, to lowest order in the perturbation, read [17, 24, 26]
@@
see also the derivation in Appendix A. Equation (8) shows that the tunneling current $I$ counts the difference between single tunneling events in the two directions, weighted by the transfered quantum $q$. In contrast, the noise, being the variance of the transmitted charge, adds such events. This structure (difference/sum of tunneling rates) was shown to extend further to all odd/even higher-order cumulants [29]. From Eq. (8), one sees that the current noise is super-Poissonian, i.e. $\mathcal{S}\geq q|I|$, which holds true even for more general out-of-equilibrium distributions [24].
In the special—but most frequently studied—case when the two subsystems have the same inverse temperature $\beta_{\text{L}}=\beta_{\text{R}}\equiv\beta$, the occupation probabilities in Eq. (6) read
$p_{n}^{(\text{L})}p_{l}^{(\text{R})}=\frac{\exp[-\beta(\epsilon_{n}^{(\text{L})}+\epsilon_{l}^{(\text{R})})]}{Z_{\text{L}}Z_{\text{R}}}.$ (9)
This simplifies the rates in Eq. (6) significantly and implies the basic detailed balance relation $\Gamma_{\rightarrow}/\Gamma_{\leftarrow}=\exp(\beta\hbar\omega)$. Applying this simple relation to the current and its zero-frequency noise given in Eq. (8) leads to the out-of-equilibrium FDT [29, 24, 26]
$\mathcal{S}=qI\coth\left(\frac{\beta\hbar\omega}{2}\right).$ (10)
It relates the tunneling current $I$ to the tunneling current noise $S$, the tunneling charge $q$, the tunneling frequency $\omega$ and the inverse temperature $\beta$. Only in the strict limit $\omega\rightarrow 0$, together with the condition $\beta_{\text{L}}=\beta_{\text{R}}$, the rates in Eq. (6) are equilibrium quantities and hence Eq. (10) reduces to the Nyquist-Johnson relation [58, 59]. However, when $\beta_{\text{L}}\neq\beta_{\text{R}}$, the relation (10) does not hold any longer [24, 57], and the out-of-equilibrium noise can have more intricate features. In the following, we focus on this temperature-biased setup and show how these features are limited by thermodynamic and energetic constraints on the tunneling rates and, as a consequence, on the tunneling current and its noise.
III Constraints on out-of-equilibrium tunneling rates
To highlight the critical role played by a sizable temperature difference, we spell out, in this Section and below, the temperature arguments explicitly. To this end, we denote by $\Gamma_{\rightarrow}(\beta_{\text{L}},\beta_{\text{R}})$ and $\Gamma_{\leftarrow}(\beta_{\text{L}},\beta_{\text{R}})$ the forward and backward out-of-equilibrium tunneling rates, where the first argument denotes the temperature subsystem L and the second the temperature of subsystem R. Without loss of generality, we also assume that, out of equilibrium, subsystem L is always hotter than subsystem R, i.e. $\beta_{\text{R}}>\beta_{\text{L}}$. We now compare the tunneling rates in the situation where the two subsystems have the same temperature, i.e., $\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}})$, with those under the desired out-of-equilibrium condition, i.e., $\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})$. In what follows, for conciseness, we refer to the equal-temperature tunneling rates $\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}})$ as equilibrium tunneling rates, even though we want to emphasize here that the dependence on the energy-transfer $\omega$, see Eq. (6), implies the full treatment of possible nonequilibrium conditions beyond a temperature bias, induced by an external agent. Differently from detailed-balance relations or fluctuation theorems relating backward and forward processes to each other but at a given nonequilibrium situation or protocol [7, 15], we here compare forward to forward and backward to backward rates, however at different nonequilibrium conditions. We show that these tunneling rates, which are transport quantities, are related to each other via inequalities that also contain thermodynamic quantities like the internal energy $U_{\alpha}(\beta)$ and the entropy $S_{\alpha}(\beta)$ for $\alpha=\mathrm{L,R}$. These are defined as
$U_{\alpha}(\beta)$ $\equiv\text{Tr}_{\alpha}\left\{\hat{H}_{\alpha}\frac{e^{-\beta\hat{H}_{\alpha}}}{Z_{\alpha}(\beta)}\right\},$ (11)
$S_{\alpha}(\beta)$ $\equiv-\text{Tr}_{\alpha}\left\{\frac{e^{-\beta\hat{H}_{\alpha}}}{Z_{\alpha}(\beta)}\log\frac{e^{-\beta\hat{H}_{\alpha}}}{Z_{\alpha}(\beta)}\right\}.$ (12)
Below, we present two constraints on the tunneling rates: a thermodynamic and an energetic one. We further discuss how these constraints are relevant for small-scale systems which do not approach the thermodynamic limit.
III.1 Thermodynamic constraint
Figure 1: Thermodynamic cycle of a fictitious tunneling experiment. (a) The two subsystems have the same (cold) temperature $T_{\text{R}}=\beta_{\text{R}}^{-1}$, and tunneling events are described by the equilibrium rates $\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}})$. (b) The left-hand (L) side is put into contact with a hot bath and heats up. In this process, the extracted work is at most $-\Delta F_{\text{L}}^{\text{(h)}}$ [Eq. (13)]. (c) The subsystems have now different temperatures, and tunneling events are described by the out-of-equilibrium rates $\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})$. (d) The system is brought back to the initial condition by putting the L side in contact with a cold bath. In this cooling process, the extracted work is at most $-\Delta F_{\text{L}}^{\text{(c)}}$ [Eq. (14)].
To introduce the thermodynamic constraint, we consider the thermodynamic cycle in Fig. 1 that a fictitious experiment could undergo. The setup consists of the two subsystems L and R, between which the tunneling transitions occur, and two macroscopic baths fixed at the inverse temperatures $\beta_{\text{L}},\beta_{\text{R}}$. These baths are used to change the temperature of the L subsystem. Initially, the fictitious experiment is prepared with the subsystems in equilibrium at the same cold temperature $T_{\text{R}}=\beta_{\text{R}}^{-1}$. Then, the following protocol is imagined to be carried out:
  • (a)
    Equilibrium rates: Tunneling events occur between L and R, according to the rates $\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}})$, i.e., in the situation where both subsystems are at the cold temperature $T_{\text{R}}$.
  • (b)
    Heating: The hot bath at temperature $T_{\text{L}}=\beta_{\text{L}}^{-1}$ is brought into contact with subsystem L. Then, heat flows from the bath into subsystem L until the latter is heated up to $T_{\text{L}}$. This process induces a change in subsystem L’s internal energy, given as $U_{\text{L}}(\beta_{\text{L}})-U_{\text{L}}(\beta_{\text{R}})\equiv\Delta U_{\text{L}}$ and of its entropy $S_{\text{L}}(\beta_{\text{L}})-S_{\text{L}}(\beta_{\text{R}})\equiv\Delta S_{\text{L}}$. Here, $U_{\text{L}}$ and $S_{\text{L}}$ are given in Eqs. (11) and (12), respectively. While heat flows, one may extract work, which is at most $-\Delta F_{\text{L}}^{\text{(h)}}$ (see Appendix C for additional details), where
    $\Delta F_{\text{L}}^{\text{(h)}}\equiv\Delta U_{\text{L}}-T_{\text{L}}\Delta S_{\text{L}},$ (13)
    is the nonequilibrium free energy change. After the heating process, we have reached the out-of-equilibrium condition of interest, where the two subsystems have different temperatures.
  • (c)
    Out-of-equilibrium rates: The cold bath is now disconnected from subsystem L, and tunneling occurs between the subsystems at the out-of-equilibrium tunneling rates $\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})$.
  • (d)
    Cooling: In this step, the total system is returned to the initial condition. This is done by bringing the cold bath at temperature $T_{\text{R}}$ in contact with subsystem L. Then, heat flows out of subsystem L until it reaches temperature $T_{\text{R}}$, thus inducing the change $-\Delta U_{\text{L}}$ and $-\Delta S_{\text{L}}$ in internal energy and entropy, respectively. Similarly to the heating step (b), while heat is flowing it is possible to extract work, which is at most $-\Delta F_{L}^{\text{(c)}}$, where
    $\Delta F_{\text{L}}^{(\text{c})}\equiv-\Delta U_{\text{L}}+T_{\text{R}}\Delta S_{\text{L}}$ (14)
    is the nonequilibrium free energy change.
With the quantities encountered during the cycle (a)-(d), we next define the thermodynamic cost function
$\begin{split}\mathcal{W}^{\text{(Thermo)}}_{\rightleftarrows}&\equiv-\Delta F_{\text{L}}^{\text{(c)}}\eta^{\text{(h)}}\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})+\\ &\qquad-\Delta F_{\text{L}}^{\text{(h)}}\eta^{\text{(c)}}\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}}),\end{split}$ (15)
where $\Delta F_{\text{L}}^{\text{(h,c)}}$ is given in Eqs. (13, 14) and $\eta^{\text{(c)}}\equiv\frac{\beta_{\text{L}}}{\beta_{\text{R}}-\beta_{\text{L}}}=\frac{T_{\text{R}}}{T_{\text{L}}-T_{\text{R}}}$ and $\eta^{\text{(h)}}\equiv\frac{\beta_{\text{R}}}{\beta_{\text{R}}-\beta_{\text{L}}}=\frac{T_{\text{L}}}{T_{\text{L}}-T_{\text{R}}}$ correspond to the coefficient of performance of a refrigerator and of a heat pump, respectively. Crucially, the product $-\Delta F_{\text{L}}^{\text{(h)}}\eta^{\text{(c)}}\equiv\widetilde{Q}^{\text{(h)}}$ sets a lower limit on the heat absorbed by the cold L subsystem during the heating process when the extracted work is maximum, as detailed in Appendix C. Similarly, the product $-\Delta F_{\text{L}}^{\text{(c)}}\eta^{\text{(h)}}\equiv\widetilde{Q}^{\text{(c)}}$ sets a lower limit on the heat absorbed by the cold bath during the cooling process. Thus, in the thermodynamic cost function (15), the equilibrium and out-of-equilibrium tunneling rates $\Gamma_{\rightleftarrows}$ are weighted by the heat dissipated to reach the opposite, namely out-of-equilibrium and equilibrium, conditions.
The thermodynamic cost function satisfies the following inequality (see Appendix D for the derivation)
$\mathcal{W}^{\text{(Thermo)}}_{\rightleftarrows}\geq\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows},$ (16)
where the rate response $\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows}$ is defined as
$\mathcal{W}^{\mathrm{(Resp)}}_{\rightleftarrows}\equiv\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})-\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}}),$ (17)
and $\partial_{\text{L}}\Gamma_{\rightleftarrows}(x,y)\equiv\frac{\partial\Gamma_{\rightleftarrows}}{\partial x}(x,y)$ is the derivative with respect to the first argument of $\Gamma_{\rightleftarrows}(x,y)$. The rate response $\mathcal{W}^{\mathrm{(Resp)}}_{\rightleftarrows}$ accounts for both the equilibrium and out-of-equilibrium response of the tunneling rates to a change in the temperature of subsystem L. Equation (16) is the first central result of this work. It implies that the response of the (nonequilibrium) rates to a change in temperature is limited by the rates themselves and by the thermodynamic cost for setting the temperature bias.
We emphasize that Eq. (16) directly translates into a constraint on the out-of-equilibrium noise, which is given by the sum of the tunneling rates via Eq. (8b). This also means, from a more practical, experimental viewpoint, that one possibility to test the constraint (16) is by measuring the noise and tunneling current in two configurations: (i) When both subsystems have the same cold temperature, and (ii) in the desired out-of-equilibrium condition. One can then access the tunneling rates as $2q^{2}\Gamma_{\rightleftarrows}=\mathcal{S}\pm qI$. All the other quantities entering the inequality can be calculated once the state of the subsystem considered at the different temperatures is known  [^1] Note that it is not needed to know the Hamiltonian of the subsystem kept at fixed temperature, which may be arbitrarily complicated, to test these constraints..
From the thermodynamic constraint (16), we can derive a direct lower bound on the out-of-equilibrium tunneling rates $\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})$ in terms of the equilibrium rates (see the derivation in Appendix E). These constraints furthermore contain the rate response, (integrals of) thermodynamic quantities and it reads
@@ III.1 Thermodynamic constraint
where we define
$\begin{split}f(x)&\equiv\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}})-\Delta F_{\text{L}}^{\text{(c)}}(x)\eta^{\text{(h)}}(x)\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}}),\\ g(x)&\equiv-\Delta F_{\text{L}}^{\text{(h)}}(x)\eta^{\text{(c)}}(x).\end{split}$ (19)
While the bound (18) has a more complex shape, containing integrals over thermodynamic functions, it has the important advantage that it does not depend on the out-of-equilibrium responses $\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})$. Instead, only the more easily accessible equilibrium response function, $\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}})$, enters Eq. (18). We emphasize that Eq. (18) directly provides a lower bound for the out-of-equilibrium noise. In particular, compared to [24], it provides a nontrivial constraint on how much super-Poissonian the noise is.
III.2 Energetic constraint
Figure 2: Heating stroke of the left (L) subsystem. As the temperature of L increases from $T_{\text{R}}$ to $T_{\text{L}}$ (top to bottom panels), the energy absorbed, $dU_{\text{L}}$, and the tunneling rates, $\Gamma_{\rightleftarrows}(\beta,\beta_{\text{R}})$, are monitored at each intermediate temperature.
In this subsection, we show that the rate response functions, $\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows}$ in Eq. (17), impose another constraint on the rates $\Gamma_{\rightleftarrows}$ themselves, which, unlike the thermodynamic constraint of Eq. (16), does not focus on the thermodynamic cost required to generate the temperature bias, but rather on the energetic cost. Therefore, this second constraint does not require considering the whole cycle (a)-(d), but only the heating stroke (d).
Starting from subsystem L being at the cold temperature $T_{\text{R}}=\beta_{\text{R}}^{-1}$, we imagine slowly increasing its temperature and identifying the tunneling rates at each temperature. This procedure is depicted in Fig. 2. Next, by weighting the tunneling rates with the corresponding variation of the internal energy needed to increase the temperature of subsystem L, we identify the following energetic cost function $\mathcal{W}^{\text{(Energy)}}_{\rightleftarrows}$ and the corresponding constraint (see Appendix F for the derivation)
@@ III.2 Energetic constraint
$\mathcal{W}^{\text{(Energy)}}_{\rightleftarrows}\equiv\int_{U_{\text{L}}(\beta_{\text{R}})}^{U_{\text{L}}(\beta_{\text{L}})}\!\!\!\!\Gamma_{\rightleftarrows}(x,\beta_{\text{R}})d[U_{\text{L}}(x)]\geq\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows}.$ (20)
Equation (20) is the second central result of this work. In contrast to the thermodynamic constraint (16), Eq. (20) involves only the internal energy of subsystem L, i.e., $U_{\text{L}}(\beta)$. This difference between Eq. (16) and Eq. (20) is also reflected on the conditions required to saturate the two constraints. By construction, both thermodynamic and energetic bounds become equalities in the zero temperature-bias regime, $\beta_{\text{L}}=\beta_{\text{R}}$. However, in this regime Eqs. (1620) become trivial since $\mathcal{W}^{\text{(Thermo)}}=\mathcal{W}^{\text{(Energy)}}=\mathcal{W}^{\text{(Resp)}}=0$. Nonetheless, the constraints can be saturated non-trivially in the presence of a temperature bias when the tunneling transitions involve states of subsystem L with a specific energy, only. While for the thermodynamic constraint this energy corresponds to the energy at which the Gibbs distribution at different temperatures cross $p_{n}^{\text{(L)}}(\beta_{\text{L}})=p_{n}^{\text{(L)}}(\beta_{\text{R}})$ (see Appendix D), for the energetic constraint this energy is the internal energy of the L subsystem $U_{\text{L}}$ (see Appendix F).

### v2

@@ III.1 Weak tunnel coupling
We study a bipartite system with Hamiltonian $\hat{H}_{0}=\hat{H}_{\text{L}}+\hat{H}_{\text{R}}$, where subsystems L (left) and R (right) may be taken as generic systems, possibly with strong interactions. Here, we take subsystem L as the system on which measurements are performed and subsystem R as the one inducing transition in subsystem L; this allows us to directly apply our results from Sec. II. The subsystems are coupled to each other by the tunneling Hamiltonian
@@
which we assume to be a small perturbation, and which induces transitions in the subsystems. Further, in the weak tunnel-coupling regime, we describe the total system with the product state $\hat{\rho}=\hat{\rho}^{\mathrm{L}}\otimes\hat{\rho}^{\mathrm{R}}$, where $\hat{\rho}^{\alpha}$ is the Gibbs state, defined in Eq. (3), at inverse temperature $\beta_{\alpha}$. With this, the rates $\Gamma_{\rightleftarrows}$ for absorbing or emitting a quantum of energy $\hbar\omega$, induced by the coupling, are—in the spectral representation [18]—given as
@@
$\begin{split}\Gamma_{\rightarrow}&\equiv\sum_{nm}\left[\frac{2\pi}{\hbar}\sum_{lk}|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon^{\text{(L)}}_{mn}+\epsilon^{\text{(R)}}_{kl}-\hbar\omega)p^{\text{(R)}}_{l}\right]p^{\text{(L)}}_{n},\\ \Gamma_{\leftarrow}&\equiv\sum_{nm}\left[\frac{2\pi}{\hbar}\sum_{lk}|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon^{\text{(L)}}_{mn}+\epsilon^{\text{(R)}}_{kl}-\hbar\omega)p^{\text{(R)}}_{k}\right]p^{\text{(L)}}_{m}.\end{split}$ (21)
See Appendix C for the derivation starting from two-point measurement transition probabilities. Here, $\mathcal{A}_{nmlk}\equiv\braket{mk}{\hat{A}}{nl}$ is the matrix element of $\hat{A}$ in the basis of the eigenstates of the generic Hamiltonian $\hat{H}_{0}$, namely $\hat{H}_{\text{L}}\ket{nl}=\epsilon^{\text{(L)}}_{n}\ket{nl},$ and $\hat{H}_{\text{R}}\ket{nl}=\epsilon^{\text{(R)}}_{l}\ket{nl}$. Furthermore, we have defined the energy differences $\epsilon^{(\alpha)}_{mn}\equiv\epsilon^{(\alpha)}_{m}-\epsilon^{(\alpha)}_{n}$, and the occupation probability of the state |$\ket{n}$ as $p^{(\alpha)}_{n}\ket{n}=\hat{\rho}^{\alpha}\ket{n}$. The terms in the square brackets in Eq. (21), which are multiplied by the initial probabilities $p_{n/m}^{(\mathrm{L})}$, represent the conditional probabilities per unit time for system L to transition from a given state $n$ to state $m$ ($\Gamma_{\rightarrow}$) with absorption of $\hbar\omega$ or from $m$ to $n$ ($\Gamma_{\leftarrow}$) with emission of $\hbar\omega$. We can thus, as presented in Sec. III.2 below, use the developed constraints for the transition rates $\Gamma_{\leftrightarrows}$, which are sums over two-point measurement transition rates. Note that the partition of the system into two weakly coupled subsystems is obviously not required for the general constraints of Eqs. (12) and (17), but importantly, it here allows us to implement a well-defined and meaningfull temperature bias. Indeed, the tunneling rates depend on two temperatures, $\Gamma_{\leftrightarrows}\equiv\Gamma_{\leftrightarrows}(\beta_{\text{L}},\beta_{\text{R}})$, via the occupation probabilities $p_{n}^{(\mathrm{L})}$ and $p_{l}^{(\mathrm{R})}$, see Eq. (21).
Interestingly, the transition rates in the weak tunneling regime can be directly connected to transport quantities, namely to a current and its zero-frequency noise [18, 25]
@@
where $q$ is a generalized charge, defined in terms of an operator $\hat{Q}$ satisfying $[\hat{Q},\hat{H}_{0}]=0$ and $[\hat{Q},\hat{A}]=q\hat{A}$, see the derivation in Appendix D.
III.2 Constraints in the tunneling regime
The thermodynamic and energetic constraints on the nonequilibrium tunneling rates—in the presence of a temperature bias—can be written starting from Eqs. (12) and (17),
$\mathcal{W}^{\text{(Thermo)}}_{\rightleftarrows}$$\geq$ $\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows},$ (23a)
$\mathcal{W}^{\text{(Energy)}}_{\rightleftarrows}$$\geq$ $\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows},$ (23b)
@@ III.2 Constraints in the tunneling regime
$\mathcal{W}^{\mathrm{(Resp)}}_{\rightleftarrows}$$\equiv$ $\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})-\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}}),$ (24a)
$\mathcal{W}^{\text{(Thermo)}}_{\rightleftarrows}$$\equiv$ $-\Delta F_{\text{L}}^{\text{(c)}}\eta^{\text{(h)}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})+$ (24b)
$-\Delta F_{\text{L}}^{\text{(h)}}\eta^{\text{(c)}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}}),$
$\mathcal{W}^{\text{(Energy)}}_{\rightleftarrows}$$\equiv$ $\int_{U_{\text{L}}(\beta_{\mathrm{c}})}^{U_{\text{L}}(\beta_{\mathrm{h}})}\!\!\!\!\Gamma_{\rightleftarrows}(x,\beta_{\mathrm{c}})d[U_{\text{L}}(x)]$ (24c)
Importantly, the rates here depend on two temperatures. We write this temperature dependence out explicitly, where the first argument is always the temperature of the left system and the second one the temperature of the right system. Note that we could have chosen any temperature $T_{\text{R}}=\beta_{\text{R}}^{-1}$, as becomes clear from the derivation of the general constraints on transition probabilities in Sec. II. Here, however, we choose one of the settings in the absence of a temperature bias, where the two subsystems have equal temperatures, in order to be able to compare to an easily accessible, experimentally relevant reference situation when establishing constraints for the out-of-equilibrium tunneling rate in the presence of a temperature bias. We therefore deliberately choose $\beta_{\text{R}}\equiv\beta_{\mathrm{c}}$ and hence compare the tunneling rates when the two subsystems have the same temperature, i.e., $\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}})$, with those under the desired out-of-equilibrium condition, i.e., $\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})$. The rate response $\mathcal{W}^{\mathrm{(Resp)}}_{\rightleftarrows}$ accounts for both the equilibrium and out-of-equilibrium response of the tunneling rates to a change in the temperature of subsystem L. Derivatives with respect to the first temperature argument are indicated in the cost function for the temperature response by $\partial_{\mathrm{L}}$. In what follows, for conciseness, we refer to the equal-temperature tunneling rates $\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}})$ as equilibrium tunneling rates, even though we want to emphasize here that the dependence on the energy transfer $\hbar\omega$, see Eq. (21), implies the full treatment of possible nonequilibrium conditions beyond a temperature bias, induced by an external agent.